How fast must an electron move so that its total e
Answer #1

The rest energy of the electron is

Eo= moc2

We calculate energy

In this case we must include the relativistic correction

E= m c2/ sqrt ( 1- (u/c)2)

They tell us that energy is 17% = 0.17 greater than the rest energy

E = Eo + 0.17 Eo = 1.17 Eo

moc2/ sqrt ( 1- (u/c)2) = 1.17 moc2

1/sqrt ( 1- (u/c)2) = 1.17

1 = 1.172( 1- (u/c)2)

1/1.172 = 1 – (u/c)2

(u/c)2= 1 – 0.7305 = 0,2695

u = c sqrt (0.2695)

u = 0.519 c